solve for X
5X + 3Y = 8X - 9Y

Found 2 solutions by CharlesG2, richwmiller:
Answer by CharlesG2(834)

solve for X
5X + 3Y = 8X - 9Y
put x's on one side and y's on other
5X-8X=-9Y-3Y --> -3X=-12Y
X=(-12/-3)Y=4Y
X=4Y

Answer by richwmiller(17219)

To solve for x we want just x on one side.
Subtract 5x from both sides
Add 9y to both sides
Divide both sides by 3
You should have something =x
Simplify the fraction